OpenStudy (helpmeee):

What is the y-value of the vertex of the function f(x) = –(x – 3)(x + 11)? The y-value of the vertex is ?

OpenStudy (mathmate):

@HelpMEEE Do you still need a response to this question?

OpenStudy (helpmeee):

yes please

OpenStudy (mathmate):

These (including your other questions) relate to properties of parabolas. Have you studied the properties of parabolas?

OpenStudy (mathmate):

In the case of a parabola equation in its vertex form: y=a(x-h)^2+k, then (h,k) is the vertex of the parabola, AND vice versa. If a>0, it opens up, and if a<0, the parabola opens down. As you probably already know, the line of symmetry of a parabola passes through the vertex. For example, y=2(x-2)+7, the vertex is at (2,7), and the (vertical) line of symmetry passes through (2,7), which makes the line of symmetry x=2. Next, since a parabola is symmetric about its vertex, the zeroes of a parabola are equidistant from the vertex. This means that in an equation (not in vertex form) in factored form: y=(x-4)(x-8) the zeroes are at (x-4)=0, or x=4, similarly x=8. Since {4,8} are the zeroes of the parabola, the line of symmetry is x=(4+8)/2=6 due to symmetry. The y-value of the vertex can be obtained by putting x=6 into the parabola equation, namely y=(6-4)(6-8)=2(-2)=-4.